# NPLQCD collaboration

### v1

@@ Abstract
Understanding the behavior of dense hadronic matter is a central goal in nuclear physics as it governs the nature and dynamics of astrophysical objects such as supernovae and neutron stars. Because of the non-perturbative nature of quantum chromodynamics (QCD), little is known rigorously about hadronic matter in these extreme conditions. Here, lattice QCD calculations are used to compute thermodynamic quantities and the equation of state of QCD over a wide range of isospin chemical potentials. Agreement is seen with chiral perturbation theory predictions when the chemical potential is small. Comparison to perturbative QCD calculations at large chemical potential allows for an estimate of the gap in the superconducting phase, and this quantity is seen to agree with perturbative determinations. Since the partition function for an isospin chemical potential, $\mu_{I}$, bounds the partition function for a baryon chemical potential $\mu_{B}=3/2\mu_{I}$, these calculations also provide rigorous non-perturbative QCD bounds on the nuclear equation of state over a wide range of baryon densities for the first time.
preprint: MIT-CTP/5729
@@
The determination of the internal structure of neutron stars presents a long-standing and important challenge for nuclear theory. Since neutron stars were first predicted in the 1930s and observed 30 years later, many models for the structure of their interiors have been proposed, including various phases of nuclear matter, mesonic condensates and hyperonic matter, and deconfined quark cores Ozel:2016oaf ; Oertel:2016bki ; Baym:2017whm ; Mannarelli:2019hgn ; Burgio:2021vgk ; Lattimer:2021emm . As observational data and terrestrial probes of the relevant nuclear densities are not sufficiently constraining, most of these possibilities for the neutron star equation of state (EoS) remain viable. From a theoretical perspective, it is expected that neutron star interiors can be described by the Standard Model of particle physics, however in a regime where the strong interactions are non-perturbative. The numerical technique of lattice quantum chromodynamics (LQCD) is applicable at such large couplings, but is beset by a notorious sign problem at nonzero baryon chemical potential deForcrand:2009zkb , prohibiting its direct application. Consequently, theoretical approaches to the nuclear EoS are based on models and interpolations between phenomenological constraints from nuclear structure and perturbative QCD (pQCD) calculations that are valid at asymptotically large chemical potentials (see for example, Refs. Annala:2023cwx ; Semposki:2024vnp ; Han:2022rug ; Jiang:2022tps ; Raaijmakers:2019dks ; Huth:2020ozf ). In light of this, any rigorous information that can impact such analyses is of paramount importance.
@@
In this work, the first non-perturbative QCD constraint on the nuclear EoS with complete quantification of systematic uncertainties is presented. These calculations build upon the proof-of-principle, single lattice spacing results of Ref. Abbott:2023coj with improved methodology, increased statistical precision, an extrapolation to the continuum limit, and an interpolation to the physical quark masses, enabling a systematically-controlled result to be achieved for the first time. The pressure and other thermodynamic properties of low-temperature isospin-dense matter are determined over a wide range of densities and chemical potentials, spanning all scales from hadronic to perturbatively-coupled. At small values of the isospin chemical potential, $\mu_{I}$, the results are found to agree with chiral perturbation theory ($\chi$PT) son_qcd_2001 ; Kogut:2001id at next-to-leading order (NLO) Adhikari:2019zaj . At large $\mu_{I}$, the results are seen to agree with pQCD with pairing contributions Fujimoto:2023mvc . The comparison of next-to-next-to-leading-order (NNLO) pQCD predictions Freedman:1976dm ; Freedman:1976ub ; Freedman:1976xs ; Baluni:1977ms ; Kurkela:2009gj ; Fujimoto:2023mvc for the pressure with the continuum limit of the LQCD calculations provides a determination of the superconducting gap as a function of $\mu_{I}$. This is seen to agree with the leading-order perturbative calculation of the pairing gap Fujimoto:2023mvc , but is more precise. Additionally, a robust conclusion from this study is that the speed of sound in isospin-dense matter significantly exceeds the conformal limit of $c_{s}^{2}/c^{2}\leq 1/3$ over a wide range of $\mu_{I}$. A Bayesian model mixing approach that combines $\chi$PT, LQCD and pQCD information provides a determination of the zero-temperature EoS for isospin-dense QCD matter valid at all values of the isospin chemical potential for the first time.
@@
From simple path-integral relations Cohen:2003ut ; Cohen:2004qp ; Fujimoto:2023unl ; Moore:2023glb , the determination of the pressure in isospin-dense matter provides a non-perturbative, model-independent bound on the pressure of isospin-symmetric QCD matter at nonzero baryon chemical potential and hence on the nuclear EoS. The current results therefore provide a systematically-controlled QCD bound at all densities, and the impact on neutron star phenomenology is briefly discussed.
@@
For an observable $\mathcal{O}$ that only depends on the energy and particle number of the system, the thermodynamic expectation value of $\mathcal{O}$ can be computed as
@@
The energy density can be computed using the quantity $E_{n}/V$, while the number density can be computed from the expectation value of $n/V$, where $V$ is the volume of the system. Derivatives of observables can also be computed using
@@
which can be obtained directly by differentiating Eq. (3). This leads to the following expressions for the pressure,
@@
to determine the isospin chemical potential from the extracted energies. Other studies have added the isospin chemical potential directly to the QCD action Kogut:2002tm ; Kogut:2002zg ; Brandt:2017oyy ; Bornyakov:2021mfj ; Brandt:2022hwy . The primary advantage of the method used here in comparison to the canonical approach is that $\mu_{I}$ enters as an input to the calculation of thermodynamic quantities rather than being derived from the isospin charge of the LQCD data and therefore is not subject to statistical and systematic uncertainties.
@@
Table 1: Parameters of the gauge-field configurations used in this work. Ensembles were generated with $N_{f}=2+1$ flavors of quarks. The first column lists the label used to refer to the ensemble, $N_{\rm conf}$ is the number of configurations, and $\beta_{g}$ and $C_{\rm SW}$ refer to the gauge coupling and clover coefficient, respectively. The lattice spacing, $a$, is determined in Refs. Yoon:2016jzj ; Mondal:2021oot , while the lattice geometries are defined by the the spatial and temporal extents, $L$ and $L_{4}$, respectively. The bare light (up and down, $m_{ud}$) and strange ($m_{s}$) quark masses are given in lattice units and $M_{\pi}$ is the pion mass. The temperature $T=1/(aL_{4})$ is also shown.
@@
where $a$ is the lattice spacing and $\vartheta_{n}(t)$ and $\sigma_{n}(t)$ are the mean and standard deviation of $\log C_{n}(t)$ evaluated over bootstrap resamplings, respectively Abbott:2023coj ($N_{b}=2000$ bootstrap resamplings are used on each ensemble to assess the statistical uncertainties and address correlations). As in Ref. Abbott:2023coj , on each bootstrap the energy is given by the value of the effective mass on a randomly chosen time inside the effective mass plateau region.
Given the energies determined on each ensemble for systems of isospin charge $n\in\{1,\ldots,6144\}$, Eqs. (5) and (6) are used to determine the pressure, the energy density, and speed of sound. The action used in these calculations is perturbatively improved, so discretization effects are ${\cal O}(a^{2},g^{2}a)$. The mass dependence of quantities evaluated over the range of quark masses used in the calculations is expected to be described linearly in $m_{ud}\sim m_{\pi}^{2}$. Each quantity ${X\in\{P/P_{\rm SB},\epsilon/\epsilon_{\rm SB},c_{s}^{2}\}}$ (where $P_{\rm SB}=\epsilon_{SB}/3=\mu_{I}^{4}/32\pi^{2}$ is the pressure of a Stefan-Boltzmann gas) is fit with forms including arbitrary $\mu_{I}$ dependence and terms ${\cal O}(a^{2},a^{2}\mu_{I},a^{2}\mu_{I}^{2},(m_{\pi}^{2}-\overline{m}_{\pi}^{2}))$, where $\overline{m}_{\pi}=139$ MeV, with coefficients independent of $\mu_{I}$. The systematic uncertainty from the extrapolation is assessed by combining fits with all possible subsets of the terms above through model averaging Rinaldi:2019thf ; NPLQCD:2020ozd ; Jay:2020jkz . The systematic uncertainty and the statistical uncertainty are combined under the bootstrap procedure.
Figure 1: The pressure computed for each LQCD ensemble as well as for the continuum limit, physical mass extraction. Results from the Stefan-Boltzmann limit of a free Fermi gas and from NNLO pQCD without pairing or quark mass effects are shown for $\mu_{I}\gtrsim 1,500$ MeV.
@@
Equation of state for isospin-dense mater: The continuum-limit lattice QCD calculations presented above span isospin chemical potentials from just above the pion mass to values where pQCD appears to converge. Consequently, by combining the LQCD results with $\chi$PT and pQCD, the zero temperature EoS of isospin-dense matter can be described for all $\mu_{I}$ with uncertainties quantified using Bayesian inference. The functional dependence of each overlapping theoretical constraint on $\mu_{I}$ is modelled by a correlated Gaussian distribution. The ensemble of constraints is combined via a Gaussian Process (GP), following similar work for the nuclear EoS Semposki:2024vnp ; Melendez:2019izc ; Essick:2020flb ; Mroczek:2023zxo . Theoretical uncertainties of $\chi$PT are estimated from the difference between the NLO and LO results, and uncertainties in pQCD are assessed from scale variation over $\bar{\Lambda}\in\mu_{I}\times[0.5,2.0]$. Figure 3 shows the GP-model results for the speed of sound in comparison to the three theoretical inputs. With a complete quantification of the isospin-dense equation of state, phenomenological implications such as the existence of pion stars Brandt:2018bwq ; Stashko:2023gnn and the isospin effects that distinguishes pure neutron matter from symmetric nuclear matter STEINER2005325 ; LI2008113 can be further investigated.
Figure 3: The squared speed of sound as a function of the isospin chemical potential. The lattice QCD determination (red), pQCD determination (orange), and $\chi$PT determination (blue) are combined into the GP-model (grey) as discussed in the main text.
@@
where $A$ is the gluon field, $\mathcal{D}(\mu)\equiv D\!\!\!\!/+m-\mu_{\mathrm{q}}\gamma_{0}$ is the Dirac operator with chemical potential $\mu_{q}$, $S_{G}$ is the gauge action, and $\int_{\beta}[dA]$ indicates integration over gauge fields with period $\beta$ in the temporal direction. As first shown in Refs. Cohen:2003ut ; Cohen:2004qp , this partition function bounds the partition function of two-flavor QCD with equal chemical potentials for $u$ and $d$ quarks
@@
By the monotonicity of the logarithm, the above inequality directly translates into an inequality between the pressures of the two media as a function of the energy density. Consequently, the isospin-dense EoS bounds the nuclear EoS for symmetric nuclear matter. At large values of the quark chemical potentials, where pQCD is valid, this bound becomes tight as differences between the partition functions enter only at ${\cal O}(\alpha_{s}^{k})$ for $k\geq 3$ Moore:2023glb . This bound was explored in Ref. Fujimoto:2023unl based on the previous lattice QCD results Abbott:2023coj at a single lattice-spacing and unphysical quark masses. Here, Fig. 4 presents updated bounds based on the continuum limit lattice QCD results at the physical quark masses, $\chi$PT, and perturbative QCD through the GP-model. While the bounds from isospin-dense matter do not significantly constrain phenomenological nuclear equations of state within the uncertainties that are typically presented Annala:2023cwx ; Semposki:2024vnp ; Han:2022rug ; Jiang:2022tps ; Raaijmakers:2019dks ; Huth:2020ozf , the bounds are independent of modeling uncertainties that enter the nuclear EoS in the regions that are unconstrained by nuclear structure or pQCD calculations.
Figure 4: The bounds on the EoS of symmetric nuclear matter derived from the GP-model of the isospin-dense EoS. Also shown are the bound from causality.

### v2

@@ Abstract
Understanding the behavior of dense hadronic matter is a central goal in nuclear physics as it governs the nature and dynamics of astrophysical objects such as supernovae and neutron stars. Because of the non-perturbative nature of quantum chromodynamics (QCD), little is known rigorously about hadronic matter in these extreme conditions. Here, lattice QCD calculations are used to compute thermodynamic quantities and the equation of state of QCD over a wide range of isospin chemical potentials. Agreement is seen with chiral perturbation theory predictions when the chemical potential is small. Comparison to perturbative QCD calculations at large chemical potential allows for an estimate of the gap in the superconducting phase, and this quantity is seen to agree with perturbative determinations. Since the partition function for an isospin chemical potential, $\mu_{I}$, bounds the partition function for a baryon chemical potential $\mu_{B}=3\mu_{I}/2$, these calculations also provide rigorous non-perturbative QCD bounds on the symmetric nuclear matter equation of state over a wide range of baryon densities for the first time.
preprint: FERMILAB-PUB-24-0333-Tpreprint: MIT-CTP/5730
@@
The determination of the internal structure of neutron stars presents a long-standing and important challenge for nuclear theory. Since neutron stars were first predicted in the 1930s and observed 30 years later, many models for the structure of their interiors have been proposed, including various phases of nuclear matter, mesonic condensates and hyperonic matter, and deconfined quark cores Ozel:2016oaf ; Oertel:2016bki ; Baym:2017whm ; Mannarelli:2019hgn ; Burgio:2021vgk ; Lattimer:2021emm . As observational data and terrestrial probes of the relevant nuclear densities are not sufficiently constraining, most of these possibilities for the neutron star equation of state (EoS) remain viable. From a theoretical perspective, it is expected that neutron star interiors can be described by the Standard Model of particle physics, however in a regime where the strong interactions are non-perturbative. The numerical technique of lattice quantum chromodynamics (LQCD) is applicable at such large couplings, but is beset by a notorious sign problem at nonzero baryon chemical potential deForcrand:2009zkb , prohibiting its direct application. Consequently, theoretical approaches to the nuclear EoS are based on models and interpolations between phenomenological constraints from nuclear structure and perturbative QCD (pQCD) calculations that are valid at asymptotically large chemical potentials (see for example, Refs. Annala:2023cwx ; Semposki:2024vnp ; Han:2022rug ; Jiang:2022tps ; Raaijmakers:2019dks ; Huth:2020ozf ; vanDalen:2005sk ; Typel:2005ba ; Kolomeitsev:2004ff ; Liu:2001iz ; Leonhardt:2019fua ). In light of this, any rigorous information that can impact such analyses is of paramount importance.
@@
In this work, the first non-perturbative QCD constraint on the nuclear EoS with complete quantification of systematic uncertainties is presented. These calculations build upon the proof-of-principle, single lattice spacing results of Ref. Abbott:2023coj with improved methodology, increased statistical precision, an extrapolation to the continuum limit, and an interpolation to the physical quark masses, enabling a systematically-controlled result to be achieved for the first time. The pressure and other thermodynamic properties of low-temperature isospin-dense matter are determined over a wide range of densities and chemical potentials, spanning all scales from hadronic to perturbatively-coupled. At small values of the isospin chemical potential, $\mu_{I}$, the results are found to agree with chiral perturbation theory ($\chi$PT) son_qcd_2001 ; Kogut:2001id at next-to-leading order (NLO) Adhikari:2019zaj ; Andersen:2023ofv . At large $\mu_{I}$, the results are seen to agree with pQCD with pairing contributions Fujimoto:2023mvc . The comparison of next-to-next-to-leading-order (NNLO) pQCD predictions Freedman:1976xs ; Freedman:1976dm ; Freedman:1976ub ; Baluni:1977ms ; Kurkela:2009gj ; Fujimoto:2023mvc for the pressure (partial next-to-next-to-next-to-leading-order results are also available Gorda:2021znl ) with the continuum limit of the LQCD calculations provides a determination of the superconducting gap as a function of $\mu_{I}$. This is seen to agree with the leading-order perturbative calculation of the pairing gap Fujimoto:2023mvc , but is more precise. The speed of sound in isospin-dense matter is also seen to significantly exceed the conformal limit of $c_{s}^{2}/c^{2}\leq 1/3$ over a wide range of $\mu_{I}$. A Bayesian model mixing approach that combines $\chi$PT, LQCD and pQCD provides a determination of the zero-temperature EoS for isospin-dense QCD matter valid at all values of the isospin chemical potential for the first time.
@@
From simple path-integral relations Cohen:2003ut ; Cohen:2004qp ; Fujimoto:2023unl ; Moore:2023glb , the determination of the pressure in isospin-dense matter provides a non-perturbative, model-independent bound on the pressure of isospin-symmetric QCD matter at nonzero baryon chemical potential and hence on the nuclear EoS. The current results therefore provide a systematically-controlled QCD bound at all densities, and the impact on nuclear phenomenology is briefly discussed.
@@
For an observable $\mathcal{O}$ that only depends on the energy and isospin charge of the system, the thermodynamic expectation value of $\mathcal{O}$ can be computed as
@@
The energy density can be computed using the quantity $E_{n}/V$, while the isospin-charge density can be computed from the expectation value of $n/V$, where $V$ is the volume of the system. Derivatives of observables can also be computed using
@@
which results from directly differentiating Eq. (3). This leads to the following expressions for the pressure,
@@
to determine the isospin chemical potential from the extracted energies. Other studies have added the isospin chemical potential directly to the QCD action Kogut:2002tm ; Kogut:2002zg ; Brandt:2017oyy ; Bornyakov:2021mfj ; Brandt:2022hwy . The primary advantage of the method used here in comparison to the canonical approach is that $\mu_{I}$ enters as an input to the calculation of thermodynamic quantities rather than being derived from the isospin charge of the LQCD data and therefore is not subject to statistical and systematic uncertainties. A more detailed comparison to the approach of Ref. Abbott:2023coj is discussed in the Supplementary Material
@@
Table 1: Parameters of the gauge-field configurations used in this work. Ensembles were generated with $N_{f}=2+1$ flavors of quarks using a clover fermion action Sheikholeslami:1985ij and a tree-level improved Lüscher-Weisz gauge action Luscher:1984xn . The first column lists the label used to refer to the ensemble, $N_{\rm conf}$ is the number of configurations, and $\beta_{g}$ and $C_{\rm SW}$ refer to the gauge coupling and clover coefficient, respectively. The lattice spacing, $a$, is determined in Refs. Yoon:2016jzj ; Mondal:2021oot , while the lattice geometries are defined by the the spatial and temporal extents, $L$ and $L_{4}$, respectively. The bare light (up and down, $m_{ud}$) and strange ($m_{s}$) quark masses are given in lattice units and $m_{\pi}$ is the pion mass. The temperature $T=1/(aL_{4})$ is also shown.
@@
where $a$ is the lattice spacing, $\vartheta_{n}(t)$ and $\sigma_{n}(t)$ are the mean and standard deviation of $\log C_{n}(t)$, and the second equality is under the assumption of log-normality Abbott:2023coj . $N_{b}=2000$ bootstrap resamplings are used on each ensemble to assess the statistical uncertainties and address correlations. As in Ref. Abbott:2023coj , on each bootstrap the energy is given by the value of the effective mass on a randomly chosen time inside the effective mass plateau region.
Given the energies determined on each ensemble for systems of isospin charge $n\in\{1,\ldots,6144\}$, Eqs. (5) and (6) are used to determine the pressure, the energy density, and speed of sound. The action used in these calculations is perturbatively improved, so discretization effects are ${\cal O}(a^{2},g^{2}a)$. The mass dependence of quantities evaluated over the range of quark masses used in the calculations is expected to be described linearly in $m_{ud}\sim m_{\pi}^{2}$. Each quantity ${X\in\{P/P_{\rm SB},\epsilon/\epsilon_{\rm SB},c_{s}^{2}/c^{2}\}}$ (where $P_{\rm SB}=\epsilon_{SB}/3=\mu_{I}^{4}/32\pi^{2}$ is the pressure of a Stefan-Boltzmann gas) is fit with forms including arbitrary $\mu_{I}$ dependence and terms ${\cal O}(a^{2},a^{2}\mu_{I},a^{2}\mu_{I}^{2},(m_{\pi}^{2}-\overline{m}_{\pi}^{2}))$, where $\overline{m}_{\pi}=m_{\pi^{+}}=139.57039$ MeV ParticleDataGroup:2022pth , with coefficients independent of $\mu_{I}$. The systematic uncertainty from the extrapolation is assessed by combining fits with all possible subsets of the terms above through model averaging Rinaldi:2019thf ; NPLQCD:2020ozd ; Jay:2020jkz . The systematic uncertainty and the statistical uncertainty are combined under the bootstrap procedure. The lattice determinations of the pressure, energy density and speed of sound are shown in the Supplementary Material, along with further details.
Figure 1: The pressure computed for each LQCD ensemble as well as for the continuum limit, physical mass extraction. Results from the Stefan-Boltzmann limit of a free Fermi gas and from NNLO pQCD without pairing or quark mass effects are shown for $\mu_{I}\gtrsim 1500$ MeV.
@@
Equation of state for isospin-dense mater: The continuum-limit lattice QCD calculations presented above span isospin chemical potentials from just above the pion mass to values where pQCD appears to converge. Consequently, by combining the LQCD results with $\chi$PT and pQCD, the zero temperature EoS of isospin-dense matter can be described for all $\mu_{I}$ with uncertainties quantified using Bayesian inference. The functional dependence of each overlapping theoretical constraint on $\mu_{I}$ is modelled by a correlated Gaussian distribution. The ensemble of constraints is combined via a Gaussian Process (GP), following similar work for the nuclear EoS Semposki:2024vnp ; Melendez:2019izc ; Essick:2020flb ; Mroczek:2023zxo . Theoretical uncertainties of $\chi$PT are estimated from the difference between the NLO and LO results, and uncertainties in pQCD are assessed from scale variation over $\bar{\Lambda}\in\mu_{I}\times[0.5,2.0]$. Figure 3 shows the GP-model results for the speed of sound in comparison to the three theoretical inputs. A data file for evaluating this model accompanies this article. With a complete quantification of the isospin-dense equation of state, phenomenological implications such as the existence of pion stars Brandt:2018bwq ; Andersen:2022aig ; Stashko:2023gnn and the isospin effects that distinguishes pure neutron matter from symmetric nuclear matter STEINER2005325 ; LI2008113 can be further investigated.
Figure 3: The squared speed of sound as a function of the isospin chemical potential. The lattice QCD determination (red), pQCD determination (orange), and $\chi$PT determination (blue) are combined into the GP model (grey) as discussed in the main text. The dashed horizontal line shows the conformal limit.
@@
where $A$ is the gluon field, $\mathcal{D}(\mu)\equiv D\!\!\!\!/+m-\mu_{\mathrm{q}}\gamma_{0}$ is the Dirac operator with quark chemical potential $\mu_{q}$, $S_{G}$ is the gauge action, and $\int_{\beta}[dA]$ indicates integration over gauge fields with period $\beta$ in the temporal direction. As first shown in Refs. Cohen:2003ut ; Cohen:2004qp , this partition function bounds the partition function of two-flavor QCD with equal chemical potentials for $u$ and $d$ quarks
@@
By the monotonicity of the logarithm, the above inequality directly translates into an inequality between the pressures of the two media as a function of the energy density. Consequently, the isospin-dense EoS bounds the EoS for symmetric nuclear matter. At large values of the quark chemical potentials, where pQCD is valid, this bound becomes tight as differences between the partition functions enter only at ${\cal O}(\alpha_{s}^{k})$ for $k\geq 3$ Moore:2023glb . This bound was explored in Ref. Fujimoto:2023unl based on the previous lattice QCD results Abbott:2023coj at a single lattice-spacing and unphysical quark masses. Here, Fig. 4 presents updated bounds based on the continuum limit lattice QCD results at the physical quark masses, $\chi$PT, and perturbative QCD through the GP-model. While the bounds from isospin-dense matter do not significantly constrain phenomenological nuclear equations of state within the uncertainties that are typically presented Annala:2023cwx ; Semposki:2024vnp ; Han:2022rug ; Jiang:2022tps ; Raaijmakers:2019dks ; Huth:2020ozf ; vanDalen:2005sk ; Typel:2005ba ; Kolomeitsev:2004ff ; Liu:2001iz ; Leonhardt:2019fua , the bounds are independent of modeling uncertainties that enter the nuclear EoS in the regions that are unconstrained by nuclear structure or pQCD calculations.
[…]